# StarCast: A Secure and Spectrum-Efficient Group Communication Scheme for LEO Satellite Networks

### v1

@@ I Introduction
However, LEO satellite communication systems face several critical constraints that limit their scalability in accommodating a massive number of user accesses. First, the onboard communication resources of LEO satellites are inherently limited due to restricted spectrum availability and the finite number of beams available for simultaneous connections, particularly in high-density areas [4]. The emergence of DTC LEO satellites exacerbates the issue of limited licensed spectrum. Satellite network operators (SNOs) must share the already allocated 4G/5G bands with terrestrial mobile network operators (MNOs) under FCC/ITU regulations (Tab. I). Additionally, the scarcity of orbital slots, which require approval from the ITU/FCC, imposes further constraints on the deployment of new satellites. Furthermore, the high capital investment required for developing, launching, and maintaining constellations presents a significant barrier to scaling these networks in a timely manner to meet sudden spikes in connectivity demand [5, 6].
@@ II-C Secure Group Communication
Despite extensive research to enhance group communication efficiency, the challenge of ensuring secure communication without compromising spectrum efficiency remains. As next-generation communication systems evolve, the security and privacy of upcoming 6G networks are increasingly critical [30, 31, 32, 33, 34, 35]. Secure group communication can be achieved through centralized key management and end-to-end session encryption [36, 37, 38]. Another approach, broadcast encryption, ensures communication security and access control by encrypting the group key using each member’s public key. Various broadcast encryption schemes embed these encrypted keys in the ciphertext [13, 14, 15, 16]. Additionally, several efforts focus on improving protection scheme efficiency [39, 40, 41]. Attribute-based encryption originated as a fuzzy variant of identity-based encryption [42]. In CP-ABE schemes [17, 18, 19, 20, 21, 22, 23], a user’s private key is associated with specific attributes, and ciphertexts are encrypted according to an access policy. Decryption is only possible if the user’s attributes meet the policy requirements.
@@ IV-A Scheme Overview
[…]
@@ IV-B Workflow
[…]
@@ IV-C Protocol Details
[…]
@@
Arbitrary binary strings can be mapped to elements of the group $\mathbb{G}_{1}$ using a hash function $\mathcal{H}$. $\mathcal{H}$ receives two kinds of input: strings and numbers. It takes the form of $(\xi,\ell,t)$ or $(\rho,\ell,t)$, where $\xi$ is an arbitrary string, $\rho$ is a positive integer, $\ell\in\{1,2,3\}$ and $t\in\{1,2\}$. These two inputs are denoted, for simplicity, as $\xi\ell t$ and $0\rho\ell t$, respectively. For easy differentiation between them, append 00 at the beginning of the second element. For all $x\in S$ and $t=1,2$, where $\sigma_{x}\leftarrow_{R}\mathbb{Z}_{p}$, compute:
@@
Enc($\textit{pk},(\mathbf{M},\pi),m$): server picks $s_{1},s_{2}\leftarrow_{R}\mathbb{Z}_{p}$ and generates:
@@ IV-D Scheme Discussion
[…]
@@ V-A Total Data Rate from Legitimate Users
AWGN power spectral density $N_{0}$174 dBm/Hz
@@ V-B Prototyping and Evaluation of StarCast Latency Using Starlink LEO Satellite Communication
Currently, Starlink satellites with DTC capability only support emergency text services and are expected to support voice and IoT connections after 2025 [46]. Therefore, we developed a prototype of StarCast using the Starlink Gen 3 Standard Kit and its LEO satellite network as an example to test the latency of StarCast (see Fig. 5). We utilized a MacBook Pro (Apple M1 Pro Processor, 16 GB Memory) as an end device, connecting it to the Starlink antenna via a built-in WiFi signal. Additionally, we tested eight Amazon Web Services (AWS) virtual machines (free-tier EC2 instances with Ubuntu 22.04 Linux OS) as servers, located in Virginia, Ohio, California, Canada, London, Frankfurt, Tokyo, and Singapore, to evaluate the end-to-end latency of ciphertext communication. These servers transmitted ciphertext to our end device, which received the LEO communication link through the Starlink standard kit. To examine the impact of different environments on the communication link, we deployed our Starlink antenna in both urban and rural areas. For comparison, we evaluated the performance of Starlink communication against other methods, including T-Mobile’s 5G and LTE network services, as well as indoor WiFi, using the same AWS servers.
@@ VI Conclusion
In this paper, we introduce StarCast, a secure and efficient group encryption scheme designed to address the critical challenges of scalability, security, and spectrum efficiency in LEO satellite networks. By leveraging CP-ABE, StarCast enables fine-grained access control, simplifies key management, and supports simultaneous data transmission to multiple users within a beam. This approach provides better utilization of limited satellite spectrum resources while maintaining robust security features such as backward secrecy, revocability, and resistance to collusion attacks. Through our implementation and comparative evaluations against unicast and broadcast frameworks, we demonstrated StarCast’s capability to significantly enhance the performance and efficiency of secure satellite communication systems. StarCast provides a scalable and reliable solution for supporting the growing connectivity demands in remote and underserved regions, laying the foundation for advanced applications in IoT, disaster recovery, and military communications.

### v2

@@ I Introduction
However, LEO satellite communication systems face several critical constraints that limit their scalability in accommodating a massive number of user accesses. First, the onboard communication resources of LEO satellites are inherently limited due to restricted spectrum availability and the finite number of beams available for simultaneous connections, particularly in high-density areas [4]. The emergence of DTC LEO satellites exacerbates the issue of limited licensed spectrum [5]. Satellite network operators (SNOs) must share the already allocated 4G/5G bands with terrestrial mobile network operators (MNOs) under FCC/ITU regulations (Tab. I). Additionally, the scarcity of orbital slots, which require approval from the ITU/FCC, imposes further constraints on the deployment of new satellites. Furthermore, the high capital investment required for developing, launching, and maintaining constellations presents a significant barrier to scaling these networks in a timely manner to meet sudden spikes in connectivity demand [6, 7].
@@ II-C Secure Group Communication
Despite extensive research to enhance group communication efficiency, the challenge of ensuring secure communication without compromising spectrum efficiency remains. As next-generation communication systems evolve, the security and privacy of upcoming 6G networks are increasingly critical [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. Secure group communication can be achieved through centralized key management and end-to-end session encryption [41, 42, 43]. Another approach, broadcast encryption, ensures communication security and access control by encrypting the group key using each member’s public key. Various broadcast encryption schemes embed these encrypted keys in the ciphertext [14, 15, 16, 17]. Additionally, several efforts focus on improving protection scheme efficiency [44, 45, 46]. Attribute-based encryption originated as a fuzzy variant of identity-based encryption [47]. In CP-ABE schemes [18, 19, 20, 21, 22, 23, 24], a user’s private key is associated with specific attributes, and ciphertexts are encrypted according to an access policy. Decryption is only possible if the user’s attributes meet the policy requirements.
@@ IV-A Scheme Overview
The key idea of StarCast is to secure the group communication scheme for LEO satellite networks while preserving the inherent spectrum efficiency of group communication. At the physical layer, StarCast broadcasts encrypted data to all users covered by the same beam simultaneously. At the presentation layer, the cryptographic design ensures data confidentiality and guarantees that only legitimate users can decrypt the information. To achieve this, the server in StarCast sends the ciphertext $\mathit{c}$ for group communication, attaching it with an access policy $\mathcal{A}$. On the user side, a secret key $\mathit{sk}$ associated with a set of attributes can decrypt the ciphertext only if the set of attributes $S$ satisfies the access policy $\mathcal{A}$. Our formal definitions of the cryptographic design follow [22, 21], whose security properties and proofs are well established.
@@ IV-B Workflow
For example, we consider the server to define the access policy for users as follows: Only the team leaders or Team Blue’s members participating in task Alpha can decrypt the message, i.e., $\mathcal{A}=(\text{"Alpha"}\text{AND}(\text{"Leader"}\text{OR}(\text{"Blue"}\text{AND}\text{"Member"})))$. As a result, the user with an attribute set $S_{1}=\{\text{Alpha},\text{Blue},\\ \text{Leader},\ldots\}$ can decrypt the message, but another user with an attribute set $S_{4}=\{\text{Beta},\text{Blue},\text{Member},\ldots\}$ cannot.
@@ IV-C Protocol Details
Setup($1^{\lambda}$): On inputting a security parameter $\lambda$, the server generates $(p,\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},e,g,h)$. $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T}$ are three groups with prime order $p$, while $g$ and $h$ are two generators for $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$, respectively. Server picks $\alpha_{1},\beta_{1},\alpha_{2},\beta_{2}\leftarrow_{R}\mathbb{Z}_{p}^{*}$ and $\delta_{1},\delta_{2},\delta_{3}\leftarrow_{R}\mathbb{Z}_{p}$, computes the public key $\mathit{pk}$ and the master key $\mathit{mk}$:
@@
Arbitrary binary strings can be mapped to elements of the group $\mathbb{G}_{1}$ using a hash function $\mathcal{H}$. $\mathcal{H}$ receives two kinds of input: strings and numbers. It takes the form of $(\xi,\ell,t)$ or $(\rho,\ell,t)$, where $\xi$ is an arbitrary string, $\rho$ is a positive integer, $\ell\in\{1,2,3\}$ and $t\in\{1,2\}$. These two inputs are denoted, for simplicity, as $\xi\ell t$ and $0\rho\ell t$, respectively. For easy differentiation between them, append 0 at the beginning of the second element. For all $x\in S$ and $t=1,2$, where $\sigma_{x}\leftarrow_{R}\mathbb{Z}_{p}$, compute:
@@
[…]
@@ IV-D Scheme Discussion
StarCast can support backward secrecy and revocability, ensuring that any new user cannot decrypt the ciphertext before joining the group. To accomplish this, the master key $\mathit{mk}$ is updated before each new user joins. Additionally, the secret keys $\mathit{sk}$ for all current legitimate users are regenerated. Member revocation is realized in a similar manner by updating the $\mathit{mk}$ for all users, excluding those that need to be revoked.
Handling Dynamically Changing User Groups. StarCast is well-suited for the dynamic nature of LEO satellite communication, where the intended recipients (i.e., policies) of different messages may change dynamically. The user’s $\mathit{sk}$ is valid for decrypting any ciphertext as long as the user’s attributes satisfy the policy of that specific ciphertext. users do not need separate keys for each ciphertext; instead, they use the same attribute-based key across different policies, decrypting only those ciphertexts for which their attributes satisfy the policy. As long as a user’s attributes have not changed, there is no need to update its $\mathit{sk}$, their secret key remains the same, regardless of how many different policies are used to encrypt different messages (for example, the access policy for a new encrypted data may become more restrictive or inclusive). The ability to decrypt one ciphertext does not affect the ability to decrypt another. Each ciphertext is evaluated independently based on its policy and the recipient’s attributes.
@@ V-A Total Data Rate from Legitimate Users
AWGN power spectral density $N_{0}$ -174 dBm/Hz
@@ V-B Prototyping and Evaluation of StarCast Latency Using Starlink LEO Satellite Communication
Currently, Starlink satellites with DTC capability support only emergency text services and are expected to extend to voice and IoT after 2025 [51]. We prototyped StarCast on the Starlink Gen 3 Standard Kit and its LEO satellite network to measure latency (see Fig. 5). A MacBook Pro (Apple M1 Pro, 16 GB) served as the end device, connected to the Starlink antenna via WiFi. We also tested eight AWS virtual machines (free-tier EC2 instances, Ubuntu 22.04) located in Virginia, Ohio, California, Canada, London, Frankfurt, Tokyo, and Singapore to assess end-to-end ciphertext latency. Ciphertext was transmitted from these servers to the end device through the Starlink link. To study environmental effects, we deployed the antenna in both urban and rural areas. For comparison, we measured performance against T-Mobile 5G/LTE and indoor WiFi using the same AWS servers.
@@ VI Conclusion
In this paper, we present StarCast, a secure and efficient group encryption scheme for LEO satellite networks that addresses scalability, security, and spectrum efficiency. Leveraging CP-ABE, StarCast enables fine-grained access control, simplifies key management, and supports simultaneous transmissions to multiple users within a beam. This approach maximizes spectrum utilization while ensuring backward secrecy, revocability, and resistance to collusion. Our implementation and comparative evaluations against unicast and broadcast frameworks show that StarCast significantly improves the performance and efficiency of secure satellite communications. StarCast offers a scalable and reliable solution to meet growing connectivity demands in remote and underserved regions, laying the foundation for advanced IoT, disaster recovery, and military applications.
VII Acknowledgment
This work was supported in part by the Office of Naval Research under grants N00014-24-1-2730 and N00014-19-1-2621, the National Science Foundation under grants 2312447, 2247560, 2154929, 2332675, 2331936, and 2235232, and the Virginia Commonwealth Cyber Initiative (CCI).